If a square has a diagonal of 40 inches how long is each side of the square
Answer by stanbon(75887)

If a square has a diagonal of 40 inches how long is each side of the square
----
Let each of the sides of the square be "x":
Equation:
x^2 + x^2 = 40^2
------
2x^2 = 1600
x^2 = 800
x^2 = 400*2
x = 20*sqrt(2)
-----
Cheers,
Stan H.
=============